# Foundation Models for Brain Signals: A Critical Review of Current Progress and Future Directions

### v1

@@ 4.2 Modeling and Pretraining
Masking: The granularity of the learned representations in the EEG-FMs is influenced by the type and extent of masking used during self-supervised pretraining [124]. In models based on masked-signal reconstruction, masking is typically applied to random patches across temporal and spatial dimensions, as opposed to masking entire channels or full time segments. For example, Brant and FoME applied a $\sim$40% uniform random masking, zeroing-out the masked input signal patches across time and channel dimensions. LaBraM similarly utilized a 50% patch-level masking, while BrainWave, despite operating in the spectrogram domain, performed masking entire spectrograms of several channels. On the other hand, BrainBERT applied a more fine-grained masking strategy, masking 5% from the time domain and 5% from the frequency domain, effectively masking $\sim$10% of the input spectrogram. Mentality applied random channel masking, targeting specific channels rather than patches. NeuroLM introduced a stair-step masking scheme, in which next token is auto-regressively predicted in a channel-conditioned manner, whereas Neuro-GPT utilized a causal masking approach in the latent space, consistent with autoregressive generation.
@@ 4.3 Model Evaluations
Evaluations on common tasks: We observed that two datasets, TUAB for abnormal EEG classification (two classes) and TUEV for EEG event detection (six classes), have been widely utilized for downstream evaluation. Note that both TUAB and TUEV are scalp EEG datasets derived from the larger TUH EEG corpus. In Figure 3(a)-3(b), we compare the performances reported on some common clinical tasks. We also analyzed the impact of the pretraining data size in the x-axis. For comparison, we use AUC and F1 scores as performance metrics as they are the most commonly reported metrics for these two tasks. Linear probing results were only available in LaBraM and are indicated with a triangle in the figure. The rest of the results are based on fine-tuning experiments performed by LaBraM, NeuroLM, EEGFormer, and BIOT. We find that AUC values ranged 0.860.920.86-0.92 for TUAB abnormal EEG classification and F1-scores ranged 0.700.850.70-0.85 for TUEV event classification, highlighting a significant performance variation across models. These comparisons also suggested that the size of the pretraining data has a minimal-to-negative relation to model performance on these two tasks.

### v2

@@ 4.2 Modeling and Pretraining
Masking: The granularity of the learned representations in the EEG-FMs is influenced by the type and extent of masking used during self-supervised pretraining [124]. In models based on masked-signal reconstruction, masking is typically applied to random patches across temporal and spatial dimensions, as opposed to masking entire channels or full time segments. For example, Brant and FoME applied a $\sim$40uniform random masking, zeroing-out the masked input signal patches across time and channel dimensions. LaBraM similarly utilized a 50% patch-level masking, while BrainWave, despite operating in the spectrogram domain, performed masking entire spectrograms of several channels. On the other hand, BrainBERT applied a more fine-grained masking strategy, masking 5% from the time domain and 5% from the frequency domain, effectively masking $\sim$10of the input spectrogram. Mentality applied random channel masking, targeting specific channels rather than patches. NeuroLM introduced a stair-step masking scheme, in which next token is auto-regressively predicted in a channel-conditioned manner, whereas Neuro-GPT utilized a causal masking approach in the latent space, consistent with autoregressive generation.
@@ 4.3 Model Evaluations
Evaluations on common tasks: We observed that two datasets, TUAB for abnormal EEG classification (two classes) and TUEV for EEG event detection (six classes), have been widely utilized for downstream evaluation. Note that both TUAB and TUEV are scalp EEG datasets derived from the larger TUH EEG corpus. In Figure 3(a)-3(b), we compare the performances reported on some common clinical tasks. We also analyzed the impact of the pretraining data size in the x-axis. For comparison, we use AUC and F1 scores as performance metrics as they are the most commonly reported metrics for these two tasks. Linear probing results were only available in LaBraM and are indicated with a triangle in the figure. The rest of the results are based on fine-tuning experiments performed by LaBraM, NeuroLM, EEGFormer, and BIOT. We find that AUC values ranged $0.86-0.92$ for TUAB abnormal EEG classification and F1-scores ranged $0.70-0.85$ for TUEV event classification, highlighting a significant performance variation across models. These comparisons also suggested that the size of the pretraining data has a minimal-to-negative relation to model performance on these two tasks.
@@ 9 Funding
Supplement
Table 3: Channel-hours were used to represent and compare the data scale at which the EEG-FMs were pretrained. In most cases, channel-hours per dataset were calculated as the product of the number of channels and the recording durations (in hours) obtained from the original studies. Scalp EEG datasets had a fixed channel count across subjects, whereas iEEG datasets, used by BrainBERT and Brant, had a subject-specific count. For Neuro-GPT and EEGFormer, channel-hours were computed from publicly available dataset statistics and were assumed to be applicable to both studies. For BrainWave, the total patch count reported was used to calculate the channel-hours by converting the patch size (1 second) to hours. Channel-hours pooled across all pretraining data sources are highlighted in bold. N/A represents cases where channel-hours could not be determined.
EEG-FM Pretraining Dataset Number of Channels Total Length of Recordings (Hours) Channel-Hours
($=$ Number of Channels $\times$ Recording Length)
Total Channel-Hours
BrainBERT Brain TreeBank [80] - - 91$\times$(1.83+1.75+0.5) + 100$\times$(2.42+2.42+2.5+3+4.33+1.75+3.42) + 91$\times$(1.92+2.67+93.92) + 135$\times$(2.58+1.33+1.83) + 205$\times$(1.75+1.77) + 152$\times$(1.15+1.68+2.43) + (109$\times$1.5) + (102$\times$2.28) + (72$\times$1.28) + 173$\times$(1.58+2.17) 14,752.68
Neuro-GPT TUEG [83] 20 27,063 541,260 541,260
Brant Private dataset (unknown source) - - (124$\times$235.34)+ (52$\times$82.39)+ (120$\times$393.09)+ (133$\times$137.45)+ (116$\times$214.22)+ (101$\times$386.70) + (67$\times$111.55) + (47$\times$207.42)+ (134$\times$759.80) 281,860.10
BIOT SHHS [85, 76] 2 42,446 84,892
PREST 16 14,197 227,152 312,044
EEGFormer TUEG 20 27,063 541,260 541,260
LaBraM BCI Competition IV-1 [102] 59 8.21 484.39
Emobrain [103] 64 4.94 316.16
Grasp/Lift EEG Challenge [104] 32 11.72 375.04
Inria BCI Challenge [105] 56 29.98 1,678.88
EEG Motor Imagery Dataset [106] 64 47.3 3,027.2
Visual Categorization EEG Data [107] 64 34.35 2,198.4
Resting State EEG Data  [108] 64 3.04 194.56
SEED Series [109, 110, 111] 62 166.75 10,338.5
Siena Scalp EEG Database [84] 31 30.47 944.57
SPIS Resting State Dataset [112] 64 0.83 53.12
Brain-Invaders [113] 32 16 512
TUAR [83] 23 92.22 2,121.06
TUEP [83] 21 591.22 12,415.62
TUSZ [83] 21 1,138.53 23,909.13
TUSL [83] 23 20.59 473.57
Multiple data sources [114, 115, 116, 117, 118] 62 342.23 21,218.26 80,260.46
Mentality TUSZ N/A N/A N/A N/A
NeuroLM TUEG 21 24,000 504,000
SEED Series 62 170.54 10,573.48
BCI Competition IV-1 59 8.21 484.39
Emobrain 64 4.94 316.16
Grasp/Lift EEG Challenge 32 11.72 375.04
Inria BCI Challenge 56 29.98 1,678.88
Motor Movement/Imagery Dataset 64 47.3 3,027.2
Raw EEG Data 64 34.35 2,198.4
Resting State EEG Data 64 3.04 194.56
Siena Scalp EEG Database 31 30.47 944.57
SPIS Resting State Dataset 64 0.83 53.12
Brain-Invaders 32 16 512
Multiple data sources [114, 115, 116, 117, 118] 62 342.23 21,218.26 545,576.06
FoME Multiple datasets N/A N/A N/A N/A
BrainWave Multiple datasets - - 3,162,233,694$\times$(1$\div$3,600) 878,398.25